A 2.0kg mass moves at a velocity of 1.0ms^-1 southward. What is the constant force which causes the mass to move at a velocity of 1.0m^-1 eastward 6.0s later?

a)0N

b)0.3N in the southeast direction

c)0.33N in the northwest direction

d)0.47N in the southwest direction

e)0.47N in the northeast direction
Answer #1

Change in velocity

V = i - ( - j) = i + j

Acceleration

a = V/t

a = (i +j) / 6

Magnitude of acceleration

a = ( 1/6) sqrt ( 1 +1)

a = 0.236

Force is given by

F = ma = 2 * 0.236 = ​​​​​​0.47 N (North east direction)

Comment in case any doubt please rate my answer....